FriskIsNice:

Can someone please help? I’ll put the question in a second. Please, I need it. I can slowly feel my heart break over the yelling. Please..

FriskIsNice:

SmokeyBrown:

The perimeter of the triangle is the sum of the three side lengths. The two side lengths we have, 20 and 16, add up to 36. And we know the total perimeter is 48. So, 48 - 36 should give us our answer, the length of the remaining side

FriskIsNice:

@smokeybrown wrote:
The perimeter of the triangle is the sum of the three side lengths. The two side lengths we have, 20 and 16, add up to 36. And we know the total perimeter is 48. So, 48 - 36 should give us our answer, the length of the remaining side
Thank you, can you keep helping me? I just need to get it up to 80%

SmokeyBrown:

No prob

SmokeyBrown:

sorry to hear about your heartbreak </3

FriskIsNice:

@smokeybrown wrote:
sorry to hear about your heartbreak </3
It’s fine, I mtrying to make it up to my mother y getting it to 80%. Even though she’s not the one stressing, can you please help me a bit more till it reaches 80

SmokeyBrown:

Yeah go for it little dude

FriskIsNice:

@smokeybrown wrote:
Yeah go for it little dude
Thank you. I know this is easy I just can’t remember how to do much of this.

SmokeyBrown:

So for rectangles we know that opposite sides have the same length. That means 2 sides have length 13 and 2 sides have length 14 To get the total perimeter we can just add 13+13+14+14

SmokeyBrown:

Also you probably wanna cut your name out of the screenshot ;)

FriskIsNice:

@smokeybrown wrote:
Also you probably wanna cut your name out of the screenshot ;)
Oof sorru, I’m trying to keep track on the score.

FriskIsNice:

SmokeyBrown:

The two sides that we know have length 8 and 13, and the total perimeter is 39. 39 - 13 - 8 means there's still 18 lengths remaining of the total perimeter. Since the remaining two sides (t) are both equal we can divide 18 by 2 to find the value of t

FriskIsNice:

@smokeybrown wrote:
The two sides that we know have length 8 and 13, and the total perimeter is 39. 39 - 13 - 8 means there's still 18 lengths remaining of the total perimeter. Since the remaining two sides (t) are both equal we can divide 18 by 2 to find the value of t
Idk how many thank you’d could be enough.

SmokeyBrown:

Oh that's a nice one. You can subtract the sides you have from the total perimeter to get the remaining sides. 21.9 - 7.3 - 7.3 happens to also be 7.3, so it seems like the sides of this triangle are equal

SmokeyBrown:

No need for thanks, really. That's what the site is for. It was in the song and everything

FriskIsNice:

Okay, and how does my teachers expect us to remember this- I was last taught this in I think 4th grade, and I’m in 7th grade.

SmokeyBrown:

Oh yeah that's pretty similar to the other trapezoid problem from just now So you can take the total perimeter, 45. Then we'll subtract the sides we know, 6 and 19. 45 - 19 - 6 gives us 20 remaining lengths. Since the last two sides are equal, we know that c is equal to 20/2

FriskIsNice:

I kinda did it by myself first so I wanna make sure I’m right- I got: 43.2

SmokeyBrown:

Yup, your math checks out! 10.8 * 4 = 43.2, so that's right

FriskIsNice:

@smokeybrown wrote:
Yup, your math checks out! 10.8 * 4 = 43.2, so that's right
okau I kinda got it I just dunno how to find the missing value-

SmokeyBrown:

We could represent this one algebraically, if that helps We know the total perimeter, 34.8 is the total of all four sides. We also know that the opposite sides of a rectangle are equal to each other From that, we can say that 34.8 = (2*13.1) + (2* u)

SmokeyBrown:

We can simplify that a little more: 34.8 = 26.2 + (2*u) Then we'll subtract 26.2 from both sides of the equation: 34.8 - 26.2 = 26.2 - 26.2 + (2*u) The result of that subtraction is: 8.6 = 2*u So u is equal to 8.6/2

FriskIsNice:

Is it just me who has anxiety, and what triggers it is yelling-

SmokeyBrown:

Nah you're in pretty good company there, trust me

SmokeyBrown:

Pretty much the same concept as before for that parallelogram there 44 = (2*18.4) + (2*a) 44 = 36.8 + (2*a) 7.2 = 2*a a = 7.2/2 = 3.6

FriskIsNice:

Excuse my name on the top I’ve gotten to the point that I’m too tired to care- also when I did it I got 14-

SmokeyBrown:

It'd be best to avoid situations that trigger anxiety, of course, but that's not always predictable or possible. And no one likes getting yelled at, anyway. While it can be difficult to cope in the moment, you should try to keep in mind that just because someone is yelling at you doesn't reflect on your value as a person

SmokeyBrown:

Anyway, back to the math!

SmokeyBrown:

Another nice and neat one Like we know, opposite sides of rectangle are equal. So, we can double the lengths of the sides we're given and add them together: 5 + 5 + 2 + 2 should give us our answer

FriskIsNice:

So close to 80%! Idk if just a word problem is any harder but whatever-

SmokeyBrown:

No worries. So we know all 4 sides of a rectangle are equal. So, to find the length of one side we can just divide the total perimeter by 4 32/4 should give us our answer

FriskIsNice:

Okay and I highly believe that this is 1.

FriskIsNice:

Well- I was wrong. DIDNT go down a lot, only went down by 1 point-

SmokeyBrown:

We can take the sides we know, multiply them by 2 and add them together, like before. 2*8.7 + 2*11.9 = 17.4 + 23.8 = 41.2, should be

FriskIsNice:

SmokeyBrown:

A similar square question. Like the other one, we can multiply the length of one side by 4 to get the total perimeter 19 *4 should give us our answer. I think... 76?

FriskIsNice:

@smokeybrown wrote:
A similar square question. Like the other one, we can multiply the length of one side by 4 to get the total perimeter 19 *4 should give us our answer. I think... 76?
Sleepy mind got to me and I put 75 instead. 🤦‍♀️

FriskIsNice:

FriskIsNice:

Update:I did 2 myself and got them both right shockingly- but then I get THIS:

SmokeyBrown:

Let's start by subtracting the sides we know from the total perimeter. Two of the sides equal 15.4, so we can take 42.8 - 15.4 - 15.4 = 42.8 - 30.8 = 12 Then the remaining two sides are equal, so we can take 12/2 to get their length

FriskIsNice:

@smokeybrown wrote:
Let's start by subtracting the sides we know from the total perimeter. Two of the sides equal 15.4, so we can take 42.8 - 15.4 - 15.4 = 42.8 - 30.8 = 12 Then the remaining two sides are equal, so we can take 12/2 to get their length
I’m sorry, I was put to sleep

SmokeyBrown:

No worries, everyone needs their rest!